1/a = 1/b + 1/c
express b in term of a and c

Answer by Alan3354(69443)

1/a = 1/b + 1/c
express b in term of a and c
-----
1/a - 1/c = 1/b
(c-a)/ac = 1/b
b = ac/(c-a)